OpenStudy (anonymous):

Explain, in your own words, the step-by-step process to simplifying the expression below. Include the simplified answer in your explanation. (6x2 − 5) − (3x2 + x + 7)

OpenStudy (anonymous):

\[6x^2 - 5 - 3x^2 - x - 7 = 3x^2 - x - 12 \]

OpenStudy (anonymous):

Then you have to find the roots aplying the formula...do you know how to do it?

OpenStudy (anonymous):

ehh a little

OpenStudy (anonymous):

distribute the -1 into the second equation and then just combine the terms

OpenStudy (anonymous):

try doing it yourself and then it'll be easier to explain.

OpenStudy (anonymous):

The roots of that are really horrible...suppose the roots are A and B. And let A be negative and B positive (are all the posibilities) then you could express it like \[3x2−x−12 = (a+x)(b-x)\]

OpenStudy (anonymous):

alright jazy..cause marrion is just saying random pellet

OpenStudy (anonymous):

s....h....i....t

OpenStudy (anonymous):

dude, always stick with common terms ex: you have 8 apples, 8 pears 8 cherries 2apple+3pears-7cherries =? u cant say oh i have 22 left fruits. u have to respect the different kinds. same goes with x^3-x^2+2x-3x^3+2x^2= you have to add up or subtract or multiply or divide common terms :) get it?

OpenStudy (anonymous):

\[(6x^2 − 5) − (3x^2 + x + 7) = (6x^2 − 5) − 3x^2 - x - 7 = 6x^2 − 5 -3x^2 - x - 7 = (6x^2-3x^2) + (-7-5) - x = 3x^2 -12 - x\]